OpenStudy (anonymous):

LAST QUESTION FOR TODAY!!!!! PLEASE!!!!!! Part 1: Use complete sentences to describe a real-world scenario that could be represented by the inequality 4x + 2y >= 40. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

jimthompson5910 (jim_thompson5910):

Part 1) You have a test. There are two types of questions: multiple choice and true false. There are 4 options per multiple choice question and 2 choices per true/false question. Now say that you know that you got at least 40 points on the test. So if that's the case, and there are x multiple choice questions and y true/false questions, then this means that 4x+2y >= 40 -------------------- Part 2) The ordered pair (10,2) is a solution since 4x+2y >= 40 4(10)+2(2) >= 40 40+4 >= 40 44 >= 40 is true. This would then mean that you got 10 multiple choice and 2 true/false questions correct.